Grade 12Electric Current

A small copper sphere carrying a charge of +0.9 microcolumb is placed at a distance separated by vacuum by 3m from an other identical sphere of charge-0.9 microcolumb in vacuum. They attract each other with a certain force. The new force between them when 5*10^12 electrons are removed from each sphere is?

Prajwal
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Dear student,

Initial charge on each of the spheres=q_1=+0.9\mu C.

Final charge on first sphere when electrons are removed from them=q_2=0.9\mu C+(1.6\times 10^{-19}C)(5\times 10^{12})=1.7\mu C

Final charge on second sphere when electrons are remvod=q_3=-0.9C+(1.6\times 10^{-19}C)(5\times 10^{12})=-0.2\mu C

\implies F_{new}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q_2q_3}{d^2}=0.00034N